OpenStudy (anonymous):

lim(h→0) ((1+h)^2+1-(1^2+1))/h)

OpenStudy (anonymous):

this is the derivative of \[x^2\] at x = 1 so you should get 2

OpenStudy (experimentx):

quite not .. though answer would be 2

OpenStudy (anonymous):

\[(1+h)^2+1=1^2+2h+h^2+1=2+2h+h^2\] \[1^2+1=2\] \[2+2h+h^2-2=2h+h^2\] \[\frac{2h+h^2}{h}=\frac{h(2+h)}{h}=2+h\]

OpenStudy (experimentx):

i guess x^2+1

OpenStudy (anonymous):

yes you are right it would be the derivative of \[x^2+1\] at x = 1

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

yw